OpenStudy (anonymous):

Use the LCD method to solve the equation x+3/x+5 + x/6 = 5/6

OpenStudy (anonymous):

\[\frac{ x+3 }{ x+5 } + \frac{ x }{ 6 } = \frac{ 5 }{ 6 }\]

OpenStudy (mathstudent55):

What is the LCD of x + 5 and 6?

OpenStudy (anonymous):

6(x+5) @mathstudent55

OpenStudy (mathstudent55):

Great. Now multiply both sides of the equation by the LCD.